OpenStudy (anonymous):

Given that r cannot equal 0 and r=5w=7a, find the value of r-w in terms of a

OpenStudy (loser66):

where are you stuck?

OpenStudy (anonymous):

I don't know how to solve it

OpenStudy (loser66):

r= 7a ok?

OpenStudy (loser66):

5w = 7a divide both sides by 5 you get \(w= \dfrac{7}{5}a\) right?

OpenStudy (loser66):

now just take r -w I rewrite them: \(r =7a\\w=\dfrac{7}{5}a\\---------------\\r-w=???\)

OpenStudy (anonymous):

Thank you !

OpenStudy (loser66):

:)